### v1

@@ Abstract
Multilevel regression and poststratification (MRP) is a computationally efficient indirect estimation method that can quickly produce improved population-adjusted estimates with limited data. Recent computational advancements allow efficient, relatively simple, and quick approximate Bayesian estimation for MRP. As population health outcomes of interest including vaccination uptake are known to have spatial structure, precision may be gained by including space in the model. We test a spatial MRP model that includes term that smooth across demographics and geographic areas using a large, unrepresentative survey. We produce California county-level estimates of first-dose COVID-19 vaccination up to June 2021 using classic and spatial MRP models, and poststratify using data from the US Census Bureau’s American Community Survey. We assess validity using reported first-dose vaccination counts from the Centers for Disease Control (CDC). Neither classic nor spatial MRP models performed well, highlighting: spatial MRP may be most appropriate for richer data contexts, some demographics in the survey data are over-sampled and -aggregated, producing model over-smoothing, and a need for survey producers to share user-representative metrics that allow better benchmarking of estimates.
@@ 1 Introduction
Multi-level regression and poststratification (Gelman and Little, 1997; Little, 1993; Park et al., 2004; Wang et al., 2015; Valliant, 2020) (“classic MRP”) is a computationally efficient indirect estimation method. It is used to smooth and adjust for non-representativeness in survey samples, and can also be used as a small area estimation (SAE) method to produce subnational areal estimates where there are limited or no specific survey data available (Lopez-Martin et al., 2022). It uses partial pooling to make predictions of unknown area-level estimates using total subpopulation trends (Bafumi and Gelman, 2006), irrespective of the spatial relationships of subnational areal units. This may be concerning because many population health outcomes of interest, including disease spread or vaccination uptake, are known to have spatial structure (clear relationships across geographic space) (Waller and Carlin, 2010). Including spatial structure in the model can improve precision, though excluding space does not necessarily lead to bias. Yet, it also is well-known that health outcomes often have demographic structure and subpopulations from neighbouring places may share similarities in health and health behaviour (for example, see: Diez Roux, 2001; Diez Roux et al., 2001).
@@
Typically, spatial Bayesian methods have required computationally demanding Markov Chain Monte Carlo (MCMC) approaches that are slow to implement, especially for spatial models which are by their nature extremely complex. Now, advancements in computation using integrated nested Laplace approximations (INLA) implemented through the r-INLA package (see: www.r-inla.org, Simpson et al., 2017) allow us to approximate Bayesian inference for complex hierarchical models with and without spatial smoothing, produce estimates, and report uncertainty extremely quickly and conveniently without MCMC. In this article, we compare classic MRP and spatial MRP – in which we include a BYM2 spatial term that smooths along or within demographic categories and between areas – using INLA, and a large, unrepresentative survey: the Delphi Group at Carnegie Mellon University’s United States-specific COVID-19 Trends and Impact Survey (CTIS), administered in partnership with Facebook (Salomon et al., 2021). The CTIS provides an exciting opportunity to test spatial MRP with real-world data: Bradley et al. (2021) have shown that the CTIS overestimates state-level direct vaccination estimates by a substantial margin compared to baseline counts reported by the Centers for Disease Control and Prevention (CDC). We compare county- and aggregated state-level estimates with CDC baseline estimates, and discuss benefits and drawbacks of classic and spatial MRP, broadly.
This paper is laid out as follows: First, we discuss the specifics of the datasets used to model and validate the results (the CTIS and CDC estimates), including descriptive statistics contained in the CTIS sample. We also describe the poststratification data used (taken from US Census Bureau’s American Community survey), including the manner in which we harmonized these with the CTIS to allow for pooling. Second, we describe four Bayesian hierarchical models (two that do not account for spatial dependence, and two with spatial smoothing). We delineate the poststratification process, including the calculations used to arrive at an aggregate estimate of first-dose vaccination, and also briefly describe the spatial and non-spatial models we build and compare in our analysis. Third, we report the results of the comparison of the modelled estimates with CTIS direct survey estimates and the CDC baseline estimates at the county level, as well as county-level estimates aggregated to the state level. Finally, we discuss the implications of these results and summarize these in the conclusion.
@@ 2 Methods
This paper compares classic and spatial MRP methods by building several Bayesian hierarchical models using the United States’ specific version of the CTIS. The outcome of interest is the number of first-dose COVID-19 vaccinations as of June 30, 2021 in California counties. Demographic strata were harmonized between this survey and the poststratification data prior to modelling. To produce estimates, we modelled the probability of first-dose vaccination against COVID-19 as of June 30, 2021 at the county level by demographic strata (age and educational attainment) using the US-specific version of the CTIS. Male and female models were built separately. To produce population estimates from these, we poststratified using 2017-2021 5-year American Community Survey population estimates with the harmonized demographic strata. To assess performance, we compared model estimates, CDC reported counts, and CTIS direct survey estimates of first-dose vaccination as of June 30, 2021. Previous research suggests COVID-19 vaccine hesitancy in the US has been driven by differences in experiences of race/ethnicity, where members of historically marginalized groups were less likely to receive vaccines (e.g., Nguyen et al., 2022; Khubchandani and Macias, 2021; Willis et al., 2021). Though this trend initially held during the COVID-19 pandemic in California, there is evidence that by early 2021 COVID-19 vaccine trends were primarily driven by educational attainment: higher education was associated with greater likelihood of vaccination (Thomas et al., 2021). Trends in the CDC California first-dose vaccination estimates also suggest vaccination decisions varied by age, educational attainment, and county (race/ethnicity data were not provided) (Centers for Disease Control and Prevention, 2023a). In the absence of race/ethnicity baseline data, we focus on variables available in both the CTIS and CDC: sex, age, and educational attainment. Detailed descriptions of these data and methods, including the generalized modelling approach, are laid out below.
@@ 2.1.3 Descriptive statistics of CTIS and poststratification weights
Table 1: Direct mean survey estimates of vaccination status in the COVID-19 Trends and Symptoms Survey (CTIS) as of June 30, 2021, by sex and sge
@@
Female 18-24 years 0.540 (0.509, 0.570)
25-64 years 0.526 (0.521, 0.532)
65 years and over 0.550 (0.533, 0.566)
Male 18-24 years 0.460 (0.430, 0.491)
25-64 years 0.474 (0.468, 0.470)
65 years and over 0.450 (0.434, 0.467)
@@ 2.2.1 Model notation
Define $n_{ijk}$ as the population in age stratum $j=1,2,3=J$ (1 = 18-24 years, 2 = 25-64 years, and 3 = 65 years and over; the reason for the limited number of strata is described below in Data), and educational attainment stratum (henceforth, “education stratum”) $k=1,...,6=K$ (where 1 = less than high school, 2 = high school or equivalent, 3 = some college, 4 = associate’s degree, 5 = bachelor’s degree, and 6 = professional or graduate degree) within county $i=1,...,58=I$. For both sexes, the total population in each county is $n_{i}=\sum^{J}_{j=1}\sum^{K}_{k=1}n_{ijk}$, and the total population of California is $N=\sum^{I}_{i=1}\sum^{J}_{j=1}\sum^{K}_{k=1}n_{ijk}$. Let $y_{ijk}$ be the number of people who received a first-dose vaccination out of the total $n_{ijk}$ people in county $i$, age stratum $j$, and education stratum $k$.
@@ 2.2.2 Hierarchical models without spatial dependence
In classic MRP, geographic variation is accounted for by modelling the outcome within each independent geographic area – e.g., by assuming counties are independent and identically distributed (IID). To reflect this, we specify a non-spatial hierarchical model with $\alpha$ the intercept, fixed effects by age, $\gamma_{j}$, education, $\beta_{k}$, and an IID random effect by county, $\epsilon_{i}$,
$\begin{gathered}y_{ijk}|\theta_{ijk}\sim BetaBinomial(n_{ijk},\theta_{ijk},d),\\ logit(\theta_{ijk})=\alpha+\gamma_{j}+\beta_{k}+\epsilon_{i},\\ \epsilon_{i}\sim_{iid}Normal(0,\sigma_{\epsilon}^{2}).\end{gathered}$ (1)
with $d$ the overdispersion parameter.
It is also possible to include smoothing across demographic strata. We can specify $\beta_{k}$ as a first-order random effect (RW1) by education, with a random walk model – a form of computationally convenient Gaussian Markov Random Field (GMRF) model (Rue and Held, 2005). In this case, a RW1 random effect smooths across the adjacent neighbors. We specify a hierarchical model,
$\begin{gathered}y_{ijk}|\theta_{ijk}\sim BetaBinomial(n_{ijk},\theta_{ijk},d),\\ logit(\theta_{ijk})=\alpha+\gamma_{j}+\beta_{k}+\epsilon_{i},\\ \underline{\beta}=[\beta_{1},...,\beta_{K}]\sim RW1(\sigma^{2}_{\beta}),\\ \epsilon_{i}\sim_{iid}Normal(0,\sigma_{\epsilon}^{2}),\end{gathered}$ (2)
with a fixed effect on age, $\gamma_{j}$ and IID by county random effects $\epsilon_{i}$. To include the intercept, $\alpha$, we include a sum-to-zero constraint for identifiability required by random walk models. We also use penalized complexity (PC) prior specifications for the hyperpriors (Simpson et al., 2017).
@@ 2.2.3 Spatial models
As vaccination trends are known to have varied across Californian counties according to CDC estimates, the outcome may have spatial dependence, and some level of measurable spatial autocorrelation (Banerjee, 2016). We describe two spatial models: one that smooths across counties with based solely on an adjacency matrix of first-order (immediately geographically proximate) neighbouring counties, and one that smooths across counties’ age groups within education strata.
It may be reasonable to expect subpopulations in neighbouring counties share some commonality with respect to vaccination uptake that has not been observed in the data (i.e., a BYM2 spatial random effect that smooths unobserved variance across neighbouring counties, $\gamma_{i}$, without it also explicitly smoothing across demographics). We may thus specify a spatial hierarchical model that includes a fixed effect for age (there are too few strata for a random effect), $\lambda_{j}$, RW1 by education, $\beta_{k}$, and also shares variance between each neighbouring county using a BYM2 spatial random effect, $\gamma_{i}$:
@@
with $\alpha$ the intercept, and PC priors. As with random walk models, BYM models including BYM2 require a sum-to-zero constraint for identifiability to include the intercept, $\alpha$ (Besag et al., 1991). This represents a classic spatial Bayesian hierarchical model that aims to shrink unobserved between-county variance.
Alternatively, it may be reasonable to expect subpopulations with similar demographics in neighbouring counties may share some commonality (variance) with respect to vaccination uptake observable across demographic strata. In this model, we smooth across counties within education group. Assuming individuals within the same education group from neighbouring counties share similar probabilities of first-dose vaccination, we specify a modified form of the Besag-York-Mollié (BYM) spatial random effect (Besag et al., 1991; Besag, 1974), a BYM2 (Riebler et al., 2016) spatial random effect that includes an IID random effect on county to smooth within and between counties across education, $\gamma_{ik}$:
$\begin{gathered}y_{ijk}|\theta_{ijk}\sim BetaBinomial(n_{ijk},\theta_{ijk},d),\\ logit(\theta_{ijk})=\alpha+\lambda_{j}+\gamma_{ik},\\ \underline{\gamma_{k}}=[\gamma_{1k},...,\gamma_{Ik}]\sim BYM2(\sigma^{2}_{\gamma},\phi_{k}).\end{gathered}$ (4)
@@ 2.2.4 Multilevel regression and poststratification
To attempt to adjust for the non-representativeness of the CTIS data, we produce marginal posterior estimates of first-dose vaccination using weights built from the marginal estimates of each demographic stratum (“level”) in each county, and also stratum-specific county population totals from census data. For each model, we sample one thousand draws of the marginal posterior probabilities of first-dose vaccination within each county $i$, age $j$, and educational attainment $k$: $\theta^{(s)}_{ijk}$ where $s=1,...,S$. From this, we can take the sample’s median and 95% credibility interval: $\theta^{(s)MED}_{ijk}(\theta^{(s)LO}_{ijk},\theta^{(s)HI}_{ijk})$.
The poststratification data are taken from the 5-year 2017-2021 ACS. These data represent the population counts for each demographic stratum by age ($J=$ 3 levels), and educational attainment ($K=$ 6 levels), for each county ($I=$ 58 counties): $J\times K\times I=1,044$ rows in the poststratification table for each sex (2,088 total). Using the poststratification data, we calculate a poststratification weight, i.e., the proportion of the total population for each sex in county $i$, age $j$, and education $k$: $w_{ijk}=n_{ijk}/n_{i}$. Estimates of the proportion of the population at the state level can then be calculated: $\theta^{(s)}=\sum_{ijk}{\theta^{(s)}_{ijk}w_{ijk}}$. County-level population estimates are thus,

### v2

@@ Abstract
Multilevel regression and poststratification (MRP) is a computationally efficient estimation method that can quickly produce improved population-adjusted estimates with limited data. Recent computational advancements allow efficient, relatively simple, and quick approximate Bayesian estimation for MRP. As population health outcomes of interest including vaccination uptake are known to have spatial structure, precision may be gained by including space in the model. We test a spatial MRP model that includes terms that smooth across demographics and geographic areas using a large, unrepresentative survey. We produce California county-level estimates of first-dose COVID-19 vaccination up to June 2021 using classic and spatial MRP models, and poststratify using data from the US Census Bureau’s American Community Survey. We assess validity using reported first-dose vaccination counts from the Centers for Disease Control (CDC). Neither classic nor spatial MRP models performed well, highlighting: spatial MRP may be most appropriate for richer data contexts, some demographics in the survey data are over-sampled and -aggregated, producing model over-smoothing, and a need for survey producers to share user-representative metrics that allow better benchmarking of estimates.
@@ 1 Introduction
Multi-level regression and poststratification (Gelman and Little, 1997; Little, 1993; Park et al., 2004; Wang et al., 2015; Valliant, 2020) (“classic MRP”) is a computationally efficient estimation method. It is used to smooth and adjust for non-representativeness in survey samples, and can also be used as a small area estimation (SAE) method to produce subnational areal estimates where there are limited or no specific survey data available (Lopez-Martin et al., 2022). It uses partial pooling to make predictions of unknown area-level estimates using total subpopulation trends (Bafumi and Gelman, 2006), irrespective of the spatial relationships of subnational areal units. This may be concerning because many population health outcomes of interest, including disease spread or vaccination uptake, are known to have spatial structure (clear relationships across geographic space) (Waller and Carlin, 2010). Including spatial structure in the model can improve precision, though excluding space does not necessarily lead to bias. Yet, it also is well-known that health outcomes often have demographic structure and subpopulations from neighbouring places may share similarities in health and health behaviour (for example, see: Diez Roux, 2001; Diez Roux et al., 2001).
@@
Typically, spatial Bayesian methods have required computationally demanding Markov Chain Monte Carlo (MCMC) approaches that are slow to implement, especially for spatial models which are by their nature extremely complex. Now, advancements in computation using integrated nested Laplace approximations (INLA) implemented through the r-INLA package (see: www.r-inla.org, Simpson et al., 2017) allow us to approximate Bayesian inference for complex hierarchical models with and without spatial smoothing, produce estimates, and report uncertainty extremely quickly and conveniently without MCMC. In this article, we compare classic MRP and spatial MRP – in which we include a BYM2 spatial term that smooths along or within demographic categories and between areas – using INLA, and a large, unrepresentative survey: the Delphi Group at Carnegie Mellon University’s United States-specific COVID-19 Trends and Impact Survey (CTIS), administered in partnership with Facebook (Salomon et al., 2021). The CTIS provides an exciting opportunity to test spatial MRP with real-world data: Bradley et al. (2021) have shown that the CTIS overestimates state-level vaccination counts by a substantial margin compared to baseline counts reported by the Centers for Disease Control and Prevention (CDC). We compare county- and aggregated state-level estimates with CDC baseline estimates, and discuss benefits and drawbacks of classic and spatial MRP, broadly.
This paper is laid out as follows: First, we discuss the specifics of the datasets used to model and validate the results (the CTIS and CDC estimates), including descriptive statistics contained in the CTIS sample. We also describe the poststratification data used (taken from US Census Bureau’s American Community survey), including the manner in which we harmonized these with the CTIS to allow for pooling. Second, we describe four Bayesian hierarchical models (two that do not account for spatial dependence, and two with spatial smoothing). We delineate the poststratification process, including the calculations used to arrive at an aggregate estimate of first-dose vaccination, and also briefly describe the spatial and non-spatial models we build and compare in our analysis. Third, we report the results of the comparison of the modelled estimates with CTIS mean survey estimates and the CDC baseline estimates at the county level, as well as county-level estimates aggregated to the state level. Finally, we discuss the implications of these results and summarize these in the conclusion.
@@ 2 Methods
This paper compares classic and spatial MRP methods by building several Bayesian hierarchical models using the United States’ specific version of the CTIS. The outcome of interest is the number of first-dose COVID-19 vaccinations as of June 30, 2021 in California counties. Demographic strata were harmonized between this survey and the poststratification data prior to modelling. To produce estimates, we modelled the probability of first-dose vaccination against COVID-19 as of June 30, 2021 at the county level by demographic strata (age and educational attainment) using the US-specific version of the CTIS. Male and female models were built separately. To produce population estimates from these, we poststratified using 2017-2021 5-year American Community Survey population estimates with the harmonized demographic strata. To assess performance, we compared model estimates, CDC reported counts, and CTIS mean survey estimates of first-dose vaccination as of June 30, 2021. Previous research suggests COVID-19 vaccine hesitancy in the US has been driven by differences in experiences of race/ethnicity, where members of historically marginalized groups were less likely to receive vaccines (e.g., Nguyen et al., 2022; Khubchandani and Macias, 2021; Willis et al., 2021). Though this trend initially held during the COVID-19 pandemic in California, there is evidence that by early 2021 COVID-19 vaccine trends were primarily driven by educational attainment: higher education was associated with greater likelihood of vaccination (Thomas et al., 2021). Trends in the CDC California first-dose vaccination estimates also suggest vaccination decisions varied by age, educational attainment, and county (race/ethnicity data were not provided) (Centers for Disease Control and Prevention, 2023a). In the absence of race/ethnicity baseline data, we focus on variables available in both the CTIS and CDC: sex, age, and educational attainment. Detailed descriptions of these data and methods, including the generalized modelling approach, are laid out below.
@@ 2.1.3 Descriptive statistics of CTIS and poststratification weights
Table 1: Mean survey estimates of vaccination status in the COVID-19 Trends and Symptoms Survey (CTIS) as of June 30, 2021, by sex and age
@@
Female 18-24 years 0.54 (0.51, 0.57)
25-64 years 0.53 (0.52, 0.53)
65 years and over 0.55 (0.53, 0.57)
Male 18-24 years 0.46 (0.43, 0.49)
25-64 years 0.47 (0.47, 0.47)
65 years and over 0.45 (0.43, 0.47)
@@ 2.2.1 Model notation
Define $n_{ijk}$ as the population in age stratum $j=\{1,2,J=3\}$ (1 = 18-24 years, 2 = 25-64 years, and 3 = 65 years and over; the reason for the limited number of strata is described in section 2.1.3), and educational attainment stratum (henceforth, “education stratum”) $k=\{1,...,K=6\}$ (where 1 = less than high school, 2 = high school or equivalent, 3 = some college, 4 = associate’s degree, 5 = bachelor’s degree, and 6 = professional or graduate degree) within county $i=\{1,...,I=58\}$. Recall we are modelling male and female separately. Therefore, for each sex the total population in each county is $n_{i}=\sum^{J}_{j=1}\sum^{K}_{k=1}n_{ijk}$, and the total population of California is $N=\sum^{I}_{i=1}\sum^{J}_{j=1}\sum^{K}_{k=1}n_{ijk}$. With respect to each sex, let $y_{ijk}$ be the number of people who received a first-dose vaccination out of the total $n_{ijk}$ people in county $i$, age stratum $j$, and education stratum $k$.
@@ 2.2.2 Hierarchical models without spatial dependence
In classic MRP, geographic variation is accounted for by modelling the outcome within each independent geographic area – e.g., by assuming counties are independent and identically distributed (IID). To reflect this, we specify a non-spatial hierarchical model with $\alpha$ the intercept, fixed effects by age, $\lambda_{j}$, education, $\beta_{k}$, and an IID random effect by county, $\epsilon_{i}$,
$\begin{gathered}y_{ijk}|\theta_{ijk}\sim BetaBinomial(n_{ijk},\theta_{ijk},d),\\ logit(\theta_{ijk})=\alpha+\lambda_{j}+\beta_{k}+\epsilon_{i},\\ \epsilon_{i}\overset{iid}{\sim}Normal(0,\sigma_{\epsilon}^{2}).\end{gathered}$ (1)
with $d$ the overdispersion parameter for the beta-binomial distribution. A fixed effect is most appropriate for age because of data sparseness (i.e., too few age strata).
It is also possible to include smoothing across demographic strata. We can specify $\beta_{k}$ as a first-order random effect (RW1) by education – a form of computationally convenient Gaussian Markov Random Field (GMRF) model (Rue and Held, 2005). In this case, a RW1 random effect smooths across the adjacent neighbors. We specify a hierarchical model,
$\begin{gathered}y_{ijk}|\theta_{ijk}\sim BetaBinomial(n_{ijk},\theta_{ijk},d),\\ logit(\theta_{ijk})=\alpha+\lambda_{j}+\beta_{k}+\epsilon_{i},\\ \underline{\beta}=[\beta_{1},...,\beta_{K}]\sim RW1(\sigma^{2}_{\beta}),\\ \epsilon_{i}\overset{iid}{\sim}Normal(0,\sigma_{\epsilon}^{2}),\end{gathered}$ (2)
with a fixed effect on age $\lambda_{j}$, and IID by county random effects $\epsilon_{i}$. To include the intercept, $\alpha$, we include a sum-to-zero constraint on $\beta$ for identifiability required by random walk models (Besag et al., 1991). Additionally, we use penalized complexity (PC) prior specifications for the hyperpriors (Simpson et al., 2017). PC priors penalize deviance from a null model (in this case, one without the random effects). PC prior choices are defined by a probability statement, $P(\sigma>u)=a$, where the user specifies $u$ and $a$. Here, we specify $u=0.5$ and $a=0.1$.
@@ 2.2.3 Spatial models
As vaccination trends are known to have varied across Californian counties according to CDC estimates, the outcome may have spatial dependence, and some level of measurable spatial autocorrelation (Banerjee, 2016). We describe two spatial models: one that smooths across counties based solely on an adjacency matrix of first-order (immediately geographically proximate) neighbouring counties, and one that smooths across counties’ education strata.
It may be reasonable to expect subpopulations in neighbouring counties share some commonality with respect to vaccination uptake that has not been observed in the data (i.e., a BYM2 spatial random effect that smooths unobserved variance across neighbouring counties, $\gamma_{i}$, without it also explicitly smoothing across demographics). We may thus specify a spatial hierarchical model that includes a fixed effect for age, $\lambda_{j}$, RW1 by education, $\beta_{k}$, and also shares information between each neighbouring county using a BYM2 spatial random effect, $\gamma_{i}$ (Riebler et al., 2016) :
@@
with $\alpha$ the intercept. In this case, the BYM2 allows for a joint prior on the unstructured and structured term. As with random walk models, BYM models including BYM2 require a sum-to-zero constraint for identifiability to include the intercept, $\alpha$ (Besag et al., 1991). Again, we use PC priors where those for the RW1 model are specified as above. With respect to the PC priors on the BYM2 model, for the overall precision parameter we set $u=1$ and $a=0.01$, corresponding to a 0.99 prior probability of having residual odds ratios smaller than 2. For the mixing parameter, we set $u=0.5$ and $a=2/3$ corresponding to a 67% chance that more than 50% of the total variation of the random effect has spatial structure. This represents a classic spatial Bayesian hierarchical model that aims to shrink unobserved between-county variance (for details on the derivation of the PC prior for the BYM2 model, see Appendix 2 in Riebler et al., 2016).
Alternatively, it may be reasonable to expect subpopulations with similar demographics in neighbouring counties may share some commonality (variance) with respect to vaccination uptake observable across demographic strata. In this model, we smooth across counties within education group. Assuming individuals within the same education group from neighbouring counties share similar probabilities of first-dose vaccination, we specify a modified form of the Besag-York-Mollié (BYM) spatial random effect (Besag et al., 1991; Besag, 1974), a BYM2 spatial random effect that includes an additional IID random effect on county to smooth within and between counties across education, $\gamma_{ik}$:
$\begin{gathered}y_{ijk}|\theta_{ijk}\sim BetaBinomial(n_{ijk},\theta_{ijk},d),\\ logit(\theta_{ijk})=\alpha+\lambda_{j}+\gamma_{ik},\\ \underline{\gamma_{k}}=[\gamma_{1k},...,\gamma_{Ik}]\sim BYM2(\sigma^{2}_{\gamma k},\phi_{k}).\end{gathered}$ (4)
@@ 2.2.4 Multilevel regression and poststratification
To attempt to adjust for the non-representativeness of the CTIS data, we produce marginal posterior estimates of first-dose vaccination using weights built from the marginal estimates of each demographic stratum (“level”) in each county, and also stratum-specific county population totals from census data. For each model, we sample one thousand draws of the marginal posterior probabilities $s$ of first-dose vaccination within each county $i$, age $j$, and educational attainment $k$ cell: $\theta^{(s)}_{ijk}$ where $s=\{1,...,S\}$ samples. From this, we can take the sample’s median and 95% credibility interval: $\theta^{(s)MED}_{ijk}(\theta^{(s)LO}_{ijk},\theta^{(s)HI}_{ijk})$.
The poststratification data are taken from the 5-year 2017-2021 ACS. These data represent the population counts for each demographic stratum by age ($J=$ 3 levels), and educational attainment ($K=$ 6 levels), for each county ($I=$ 58 counties): $J\times K\times I=1,044$ rows in the poststratification table for each sex (2,088 total). Using the poststratification data, we calculate a poststratification weight, i.e., the proportion of the total population for each sex in county $i$, age $j$, and education $k$: $w_{ijk}=n_{ijk}/n_{i}$. Estimates of the proportion of the population at the state level can then be calculated: $\theta^{(s)}=\sum_{ijk}{\theta^{(s)}_{ijk}w_{ijk}}$. County-level population estimates are,